# Non-Markovian Reduced Models to Unravel Transitions in Non-equilibrium Systems

### v1

@@ II.2 Invariance equation and backward-forward systems
Before diving into the parameterization formulas retained to address these limitations (Section II.3), we recall the basics of reducing an SPDE to a Stochastic Invariant Manifold (SIM). In particular, this detour will enable us to better appreciate the emergence of BF systems as a natural framework for parameterizations. For that purpose, adapting the approach from [29], we transform the SPDE reduction problem into the more tractable problem of reducing a PDE with random coefficients. This simplification makes the problem significantly more amenable to analysis than its original SPDE form.
@@
(Ornstein-Uhlenbeck process), where ${\bm{W}_{t}}$ is defined in Eq. (4) with the corresponding mutually independent Brownian motions, $W_{t}^{j}$, in place of the $\eta_{t}^{j}$, i.e.:
@@
$\partial_{t}\phi+\mathcal{L}_{A}[\phi](X)=-D\phi(t,X,\omega)\Big{(}\Pi_{\mathfrak{c}}G(X+\phi+z(t,\omega))\Big{)}+\Pi_{\mathfrak{s}}G(X+\phi+z(t,\omega)),$ (15)
where $\mathcal{L}_{A}$ denotes the operator acting on differentiable mappings $\psi$ from $H_{\mathfrak{c}}$ into $H_{\mathfrak{s}}$, as follows:
$\mathcal{L}_{A}[\psi](X)=D\psi(X)A_{\mathfrak{c}}X-A_{\mathfrak{s}}\psi(X),\;X\in H_{\mathfrak{c}},$ (16)
with $A_{\mathfrak{c}}=\Pi_{\mathfrak{c}}A$ and $A_{\mathfrak{s}}=\Pi_{\mathfrak{s}}A$.
To simplify, assume that $H_{\mathfrak{c}}$ and $H_{\mathfrak{s}}$ are chosen such that $\Pi_{\mathfrak{s}}{\bm{W}_{t}}=0$. Then in particular $z_{\mathfrak{s}}=\Pi_{\mathfrak{s}}z=0$. Consider now the Lyapunov-Perron integral
$\mathfrak{J}(t,X,\omega)=\int_{-\infty}^{t}e^{(t-s)A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}G_{k}(e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s,\omega))\,\mathrm{d}s,$ (17)
@@
$\hskip 85.35826pt\left(\Re\Big{(}\lambda_{n}-\sum_{p\in K\setminus K_{z}}\lambda_{j_{p}}\Big{)}<0\right),$
@@
$\partial_{t}\psi+\mathcal{L}_{A}[\psi](X)=\Pi_{\mathfrak{s}}G_{k}(X+z_{\mathfrak{c}}(t,\omega)).$ (19)
@@
$\frac{\,\mathrm{d}q}{\,\mathrm{d}t}=A_{\mathfrak{s}}q+\Pi_{\mathfrak{s}}G_{k}(p+z_{\mathfrak{c}}(s,\omega)),\hskip 5.69046pts\in[t-\tau,t],$ (20b)
@@
$\lim_{\tau\rightarrow\infty}\|\mathfrak{J}(t,X,\omega)-q_{\tau}(t,X,\omega)\|_{H_{\mathfrak{s}}}=0,$ (21)
@@
Theorem II.1 teaches us that solving the BF system (20) gives the solution to the homological equation (17) which is an approximation of the full invariance equation Eq. (15). In the deterministic setting, solutions to the homological equation (17) are known to provide actual approximations of the invariant manifolds with rigorous estimates; see [27, Theorem 1] and [28, Theorem III.1]. Such results extend to the case of SPDEs driven by multiplicative (parameter) noise; see [31, Theorem 2.1] and [29, Theorem 6.1 and Corollary 7.1]. It is not the scope of this article to deal with rigorous error estimates regarding the approximation problem of stochastic invariant manifolds for SPDEs driven by additive noise, but the rationale stays the same: solutions to the homological equation (17) or equivalently to the BF system (20) provide actual approximations of the underlying stochastic invariant manifolds, in the additive noise case as well.
@@
$\,\mathrm{d}\widehat{q}=\left(A_{\mathfrak{s}}\widehat{q}+\Pi_{\mathfrak{s}}G_{k}(\widehat{p})\right)\,\mathrm{d}s,\hskip 36.98866pts\in[t-\tau,t],$ (22b)
@@ II.3 Approximations of fully coupled backward-forward systems
$\,\mathrm{d}q=\big{[}A_{\mathfrak{s}}q+\Pi_{\mathfrak{s}}G(p+q)\big{]}\mathrm{d}s+\Pi_{\mathfrak{s}}\mathrm{d}\bm{W}_{s},\;\;s\in I_{t,\tau},$ (23b)
@@
$\mathrm{d}q^{(\ell)}=\Big{[}A_{\mathfrak{s}}q^{(\ell)}+\Pi_{\mathfrak{s}}G\big{(}p^{(\ell-1)}+q^{(\ell-1)}\big{)}\Big{]}\mathrm{d}s+\Pi_{\mathfrak{s}}\mathrm{d}\bm{W}_{s},$ (25b)
@@
$q^{(1)}(s)=\epsilon\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}\,\mathrm{d}\bm{W}_{s^{\prime}},$
@@
$q^{(2)}(s)=\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}B\big{(}p^{(1)}(s^{\prime})+q^{(1)}(s^{\prime})\big{)}\,\mathrm{d}s^{\prime}$ (27)
$\hskip 56.9055pt+\epsilon\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}\,\mathrm{d}\bm{W}_{s^{\prime}}.$
@@
$\Phi(t,X,\omega)$ $=\int_{t-\tau}^{t}e^{(t-s)A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}G\big{(}e^{(s-t)A_{\mathfrak{c}}}X\big{)}\,\mathrm{d}s$ (33)
$\qquad+\epsilon\int_{t-\tau}^{t}e^{(t-s)A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}\,\mathrm{d}\bm{W}_{s}(\omega),$
restoring $B$ to be a more general nonlinearity $G$.
@@
$\!\,\mathrm{d}q=\left(A_{\mathfrak{s}}q+\Pi_{\mathfrak{s}}G(p)\right)\,\mathrm{d}s+\epsilon\Pi_{\mathfrak{s}}\,\mathrm{d}\bm{W}_{s},\;\;s\in I_{t,\tau},$ (34b)
@@
Section III.2 details an efficient computation of the single stochastic convolution in equation (33) using auxiliary ODEs with random coefficients. It is worth noting that previous SPDE reduction approaches have, in certain circumstances or under specific assumptions (e.g., $\Pi_{c}B(q,q)=0$ as seen in [38, Eq. (2.2)] and [100, Eq. (2.10)]), deliberately opted to avoid directly computing these convolutions via quadrature [101, 102, 50]. Other approaches have shown the relevance to keep Ornstein-Uhlenbeck terms such as in Eq. (33) in the approximation of low-mode amplitudes to derive rigorous error estimates; see e.g. [38, 49]. We show in Sections V and VI below that such stochastic convolutions become crucial for the reduction when the chosen cutoff scale, defining the retained large-scale dynamics, is larger than the forcing scale. Unlike previous studies, we furthermore show below that optimizing the memory length of the noise’s past ($\tau$ in Eq. (33)) is in that respect, a key ingredient. Also, as exemplified in applications, optimizing the nonlinear terms in Eq. (33) may turn to be of utmost importance to capture efficiently the average motion of the SPDE dynamics; see Section V.5 below.
@@ III Non-Markovian Parameterizations: Formulas and Optimization
where $\Pi_{n}$ denotes the projector onto the mode $\bm{e}_{n}$. Here, $\tau$ is a free parameter to be adjusted, no longer condemned to approach $\infty$ as in the approximation theory of SIM discussed above. Note that compared to Eq. (34), we break down the forcing mode by mode for each high mode (scale-awareness) to allow for adjusting the free backward parameter, $\tau$, per scale to parameterize. This strategy allows for a greater degree of freedom to calibrate useful parameterizations, as will be apparent in applications dealt with in Sections V and VI.
@@ III.1 Stochastic parameterizations for systems with cubic interactions
Here, $A$ is a partial differential operator as defined in (2) for a suitable ambient Hilbert space $H$; see also [30, Chap. 2]. The operators $G_{2}$ and $G_{3}$ are quadratic and cubic, respectively. Here again, we assume the eigenmodes of $A$ and its adjoint $A^{*}$ to form a bi-orthonormal basis of $H$; namely, the set of eigenmodes $\{\bm{e}_{j}\}$ of $A$ and $\{\bm{e}^{*}_{j}\}$ of $A^{*}$ each forms a Hilbert basis of $H$, and they satisfy the bi-orthogonality condition $\langle\bm{e}_{j},\bm{e}^{*}_{k}\rangle_{H}=\delta_{j,k}$, where $\delta_{j,k}$ denotes the Kronecker delta function. The noise term ${\bm{W}_{t}}$ takes the form defined in Eq. (12). Stochastic equations such as Eq. (36) arise in many areas of physics such as in the description of phase separation [103, 104, 45], pattern formations [3, 7] or transitions in geophysical fluid models [20, 105, 106, 55, 25, 21, 31].
@@
where the $W_{s}^{k}$ are the independent Brownian motions in (12). It is known that Eq. (36) admits a unique mild solution in $H$ with the right conditions on the coefficients of $G_{2}$ and $G_{3}$ (confining potential) [88]; see also [107, Prop. 3.4]. We refer to [89, 88] for background on the stochastic convolution term $\int_{0}^{t}e^{(t-s)A}\,\mathrm{d}\bm{W}_{s}$.
@@
$\;\;+Z^{n}_{\tau}({\bm{\lambda}},t;\omega)+\sum_{i,j=1}^{m_{c}}\Bigl{(}D^{n}_{ij}(\tau,{\bm{\lambda}})B_{ij}^{n}X_{i}X_{j}\Bigr{)}$
$\;\;+\sum_{i,j,k=1}^{m_{c}}\Big{(}E_{ijk}^{n}(\tau,{\bm{\lambda}})C_{ijk}^{n}X_{i}X_{j}X_{k}\Big{)},$
@@
$Z^{n}_{\tau}({\bm{\lambda}},t;\omega)=\sigma_{n}\lambda_{n}e^{\lambda_{n}t}\int_{t-\tau}^{t}e^{-\lambda_{n}s}W_{s}^{n}(\omega)\,\mathrm{d}s,$ (41)
@@
while the coefficients $D^{n}_{ij}(\tau,{\bm{\lambda}})$ and $E_{ijk}^{n}(\tau,{\bm{\lambda}})$ are given by
$D_{ij}^{n}(\tau,{\bm{\lambda}})=\begin{cases}\frac{1-\exp\big{(}-\delta_{ij}^{n}\tau\big{)}}{\delta_{ij}^{n}},&\text{if $\delta_{ij}^{n}\neq 0$},\\ \tau,&\text{otherwise},\end{cases}$ (43)
@@
$E_{ijk}^{n}(\tau,{\bm{\lambda}})=\begin{cases}\frac{1-\exp\big{(}-\delta_{ijk}^{n}\tau\big{)}}{\delta_{ijk}^{n}},&\text{if $\delta_{ijk}^{n}\neq 0$},\\ \tau,&\text{otherwise},\end{cases}$ (44)
@@
In what follows we often drop the dependence of $\Phi_{n}$ on ${\bm{\lambda}}$ and the noise path $\omega$. The parameterization of the neglected scales writes then as
$\Phi(\bm{\tau},X,t)=\sum_{n\geq m_{c}+1}\Phi_{n}(\tau_{n},X,t)\bm{e}_{n},\$ (45)

### v2

@@ II.2 Invariance equation and backward-forward systems
Before diving into the parameterization formulas retained to address these limitations (Section II.3), we recall the basics of reducing an SPDE to a Stochastic Invariant Manifold (SIM). In particular, this detour enables us to better appreciate the emergence of BF systems as a natural framework for parameterizations. For that purpose, adapting the approach from [30], we transform the SPDE reduction problem into the more tractable problem of reducing a PDE with random coefficients. This simplification makes the problem significantly more amenable to analysis than its original SPDE form.
@@
(Ornstein-Uhlenbeck (OU) process), where ${\bm{W}_{t}}$ is defined in Eq. (4) with the corresponding mutually independent Brownian motions, $W_{t}^{j}$, in place of the $\eta_{t}^{j}$, i.e.:
@@
$\partial_{t}\phi+\mathcal{L}_{A}[\phi](X)=-D\phi(t,X,\omega)\Big{(}\Pi_{\mathfrak{c}}G(X+\phi+z(t,\omega))\Big{)}+\Pi_{\mathfrak{f}}G(X+\phi+z(t,\omega)),$ (15)
where $\mathcal{L}_{A}$ denotes the operator acting on differentiable mappings $\psi$ from $H_{\mathfrak{c}}$ into $H_{\mathfrak{f}}$, as follows:
$\mathcal{L}_{A}[\psi](X)=D\psi(X)A_{\mathfrak{c}}X-A_{\mathfrak{f}}\psi(X),\;X\in H_{\mathfrak{c}},$ (16)
with $A_{\mathfrak{c}}=\Pi_{\mathfrak{c}}A$ and $A_{\mathfrak{f}}=\Pi_{\mathfrak{f}}A$.
To simplify, assume that $H_{\mathfrak{c}}$ and $H_{\mathfrak{f}}$ are chosen such that $\Pi_{\mathfrak{f}}{\bm{W}_{t}}=0$. Then in particular $z_{\mathfrak{f}}=\Pi_{\mathfrak{f}}z=0$. Consider now the Lyapunov-Perron integral
$\mathfrak{J}(t,X,\omega)=\int_{-\infty}^{t}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}G_{k}(e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s,\omega))\,\mathrm{d}s,$ (17)
@@
$\hskip 85.35826pt\left(\mathrm{Re}\Big{(}\lambda_{n}-\sum_{p\in K\setminus K_{z}}\lambda_{j_{p}}\Big{)}<0\right),$
@@
$\partial_{t}\psi+\mathcal{L}_{A}[\psi](X)=\Pi_{\mathfrak{f}}G_{k}(X+z_{\mathfrak{c}}(t,\omega)).$ (19)
@@
$\frac{\,\mathrm{d}q}{\,\mathrm{d}t}=A_{\mathfrak{f}}q+\Pi_{\mathfrak{f}}G_{k}(p+z_{\mathfrak{c}}(s,\omega)),\hskip 5.69046pts\in[t-\tau,t],$ (20b)
@@
$\lim_{\tau\rightarrow\infty}\|\mathfrak{J}(t,X,\omega)-q_{\tau}(t,X,\omega)\|_{H_{\mathfrak{f}}}=0,$ (21)
@@
Theorem II.1 teaches us that solving the BF system (20) gives the solution to the homological equation (17) which is an approximation of the full invariance equation Eq. (15). In the deterministic setting, solutions to the homological equation (17) are known to provide actual approximations of the invariant manifolds with rigorous estimates; see [27, Theorem 1] and [28, Theorem III.1]. Such results extend to the case of SPDEs driven by multiplicative (parameter) noise; see [31, Theorem 2.1] and [30, Theorem 6.1 and Corollary 7.1]. It is not the scope of this article to deal with rigorous error estimates regarding the approximation problem of stochastic invariant manifolds for SPDEs driven by additive noise, but the rationale stays the same: solutions to the homological equation (17) or equivalently to the BF system (20) provide actual approximations of the underlying stochastic invariant manifolds, in the additive noise case as well.
@@
$\,\mathrm{d}\widehat{q}=\left(A_{\mathfrak{f}}\widehat{q}+\Pi_{\mathfrak{f}}G_{k}(\widehat{p})\right)\,\mathrm{d}s,\hskip 36.98866pts\in[t-\tau,t],$ (22b)
@@ II.3 Approximations of fully coupled backward-forward systems
$\,\mathrm{d}q=\big{[}A_{\mathfrak{f}}q+\Pi_{\mathfrak{f}}G(p+q)\big{]}\mathrm{d}s+\Pi_{\mathfrak{f}}\mathrm{d}\bm{W}_{s},\;\;s\in I_{t,\tau},$ (23b)
@@
$\mathrm{d}q^{(\ell)}=\Big{[}A_{\mathfrak{f}}q^{(\ell)}+\Pi_{\mathfrak{f}}G\big{(}p^{(\ell-1)}+q^{(\ell-1)}\big{)}\Big{]}\mathrm{d}s+\Pi_{\mathfrak{f}}\mathrm{d}\bm{W}_{s},$ (25b)
@@
$q^{(1)}(s)=\epsilon\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}\,\mathrm{d}\bm{W}_{s^{\prime}},$
@@
$q^{(2)}(s)=\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}B\big{(}p^{(1)}(s^{\prime})+q^{(1)}(s^{\prime})\big{)}\,\mathrm{d}s^{\prime}$ (27)
$\hskip 56.9055pt+\epsilon\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}\,\mathrm{d}\bm{W}_{s^{\prime}}.$
@@
$\Phi(t,X,\omega)$ $=\int_{t-\tau}^{t}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}G\big{(}e^{(s-t)A_{\mathfrak{c}}}X\big{)}\,\mathrm{d}s$ (33)
$\qquad+\epsilon\int_{t-\tau}^{t}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}\,\mathrm{d}\bm{W}_{s}(\omega),$
after restoring $G$ as a more general nonlinearity than $B$.
@@
$\!\,\mathrm{d}q=\left(A_{\mathfrak{f}}q+\Pi_{\mathfrak{f}}G(p)\right)\,\mathrm{d}s+\epsilon\Pi_{\mathfrak{f}}\,\mathrm{d}\bm{W}_{s},\;\;s\in I_{t,\tau},$ (34b)
@@
Section III.2 details an efficient computation of the single stochastic convolution in equation (33) using auxiliary ODEs with random coefficients. It is worth noting that previous SPDE reduction approaches have, in certain circumstances or under specific assumptions (e.g., $\Pi_{c}B(q,q)=0$ as seen in [54, Eq. (2.2)] and [40, Eq. (2.10)]), deliberately opted to avoid directly computing these convolutions via quadrature [100, 101, 102].
The stochastic convolution in Eq. (33) accounts for finite-range memory effects stemming from the stochastic forcing to the original equation. Such terms have been encountered in the approximation of low-mode amplitudes, albeit in their asymptotic format when $\tau\rightarrow\infty$ see e.g. [54, 103]. As shown in our examples below (Sections V and VI), stochastic convolution terms become crucial for efficient reduction when the cutoff scale, defining the retained large-scale dynamics, is larger than the forcing scale. There, we show furthermore that the finite-range memory content (measured by $\tau$) is a key factor for a skillful reduction when optimized properly. Also, as exemplified in applications, optimizing the nonlinear terms in Eq. (33) may turn to be of utmost importance to reproduce accurately the average motion of the SPDE dynamics; see Section V.5 below.
@@ III Non-Markovian Parameterizations: Formulas and Optimization
where $\Pi_{n}$ denotes the projector onto the mode $\bm{e}_{n}$. Here, $\tau$ is a free parameter to be adjusted, no longer condemned to approach $\infty$ as in the approximation theory of SIM discussed above. Note that compared to Eq. (34), we break down the forcing mode by mode for each high mode to allow for adjusting the free backward parameter, $\tau$, per scale to parameterize (scale-awareness). This strategy allows for a greater degree of freedom to calibrate useful parameterizations, as will be apparent in applications dealt with in Sections V and VI.
@@ III.1 Stochastic parameterizations for systems with cubic interactions
Here, $A$ is a partial differential operator as defined in (2) for a suitable ambient Hilbert space $H$; see also [29, Chap. 2]. The operators $G_{2}$ and $G_{3}$ are quadratic and cubic, respectively. Here again, we assume the eigenmodes of $A$ and its adjoint $A^{*}$ to form a bi-orthonormal basis of $H$; namely, the set of eigenmodes $\{\bm{e}_{j}\}$ of $A$ and $\{\bm{e}^{*}_{j}\}$ of $A^{*}$ each forms a Hilbert basis of $H$, and they satisfy the bi-orthogonality condition $\langle\bm{e}_{j},\bm{e}^{*}_{k}\rangle_{H}=\delta_{j,k}$, where $\delta_{j,k}$ denotes the Kronecker delta function. The noise term ${\bm{W}_{t}}$ takes the form defined in Eq. (12). Stochastic equations such as Eq. (36) arise in many areas of physics such as in the description of phase separation [104, 105, 56], pattern formations [3, 7] or transitions in geophysical fluid models [20, 106, 107, 108, 25, 21, 31].
@@
where the $W_{s}^{k}$ are the independent Brownian motions in (12). It is known that Eq. (36) admits a unique mild solution in $H$ with the right conditions on the coefficients of $G_{2}$ and $G_{3}$ (confining potential) [87]; see also [109, Prop. 3.4]. We refer to [88, 87] for background on the stochastic convolution term $\int_{0}^{t}e^{(t-s)A}\,\mathrm{d}\bm{W}_{s}$.
@@
$\;\;+Z^{n}_{\tau}(t;\omega)+\sum_{i,j=1}^{m_{c}}\Bigl{(}D^{n}_{ij}(\tau)B_{ij}^{n}X_{i}X_{j}\Bigr{)}$
$\;\;+\sum_{i,j,k=1}^{m_{c}}\Big{(}E_{ijk}^{n}(\tau)C_{ijk}^{n}X_{i}X_{j}X_{k}\Big{)},$
@@
$Z^{n}_{\tau}(t;\omega)=\sigma_{n}\lambda_{n}e^{\lambda_{n}t}\int_{t-\tau}^{t}e^{-\lambda_{n}s}W_{s}^{n}(\omega)\,\mathrm{d}s,$ (41)
@@
while the coefficients $D^{n}_{ij}(\tau)$ and $E_{ijk}^{n}(\tau)$ are given by
$D_{ij}^{n}(\tau)=\begin{cases}\frac{1-\exp\big{(}-\delta_{ij}^{n}\tau\big{)}}{\delta_{ij}^{n}},&\text{if $\delta_{ij}^{n}\neq 0$},\\ \tau,&\text{otherwise},\end{cases}$ (43)
@@
$E_{ijk}^{n}(\tau)=\begin{cases}\frac{1-\exp\big{(}-\delta_{ijk}^{n}\tau\big{)}}{\delta_{ijk}^{n}},&\text{if $\delta_{ijk}^{n}\neq 0$},\\ \tau,&\text{otherwise},\end{cases}$ (44)
@@
Remark III.1
In connection with the stochastic convolution $\xi_{t,\tau}=\int_{t-\tau}^{t}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}\,\mathrm{d}\bm{W}_{s}$ in Eq. (33), we note that the stochastic terms in Eq. (40), involving the processes $W^{n}_{t}$, $W^{n}_{t-\tau}$, and $Z^{n}_{\tau}$, can be rewritten as a scalar stochastic integral, more precisely:
$\int_{t-\tau}^{t}e^{\lambda_{n}(t-s)}\,\mathrm{d}W^{n}_{t}=W^{n}_{t}-e^{\tau\lambda_{n}}W^{n}_{t-\tau}+\sigma_{n}^{-1}Z^{n}_{\tau}(t).$ (45)
Note that the RHS in Eq. (45) is simply the projection of $\xi_{t,\tau}$ onto mode $\bm{e}_{n}$ (Eq. (38)).
Equation (45) is a direct consequence of Itô’s formula applied to the product $e^{-\lambda s}W_{s}^{n}$ followed by integration over $[t-\tau,t]$. Indeed, we have (dropping the $n$-dependence),
$\,\mathrm{d}\Big{(}e^{-\lambda s}W_{s}\Big{)}$ $=-\lambda e^{-\lambda s}W_{s}\,\mathrm{d}s+e^{-\lambda s}\,\mathrm{d}W_{s}+\,\mathrm{d}(e^{-\lambda s})\,\mathrm{d}W_{s}$
$=-\lambda e^{-\lambda s}W_{s}\,\mathrm{d}s+e^{-\lambda s}\,\mathrm{d}W_{s},$
by treating $\,\mathrm{d}s\,\mathrm{d}W_{s}$ as zero due to Itô calculus [111].
By introducing now $\delta W=W_{t}-e^{\lambda\tau}W_{t-\tau}$ and denoting by $J$ the stochastic convolution $\int_{t-\tau}^{t}e^{\lambda(t-s)}\,\mathrm{d}W_{s}$, we arrive due to Eq. (41) at
$e^{-\lambda t}\delta W+\lambda e^{-\lambda t}Z_{\tau}=e^{-\lambda t}J,$ (46)
which gives the relation $J=\delta W+\lambda Z_{\tau}$, where we have set $\sigma_{n}=1$ to simplify.
In the expression of $\Phi_{n}$ given by Eq. (40), we opt for the expression involving the Riemann–Stieltjes integral $Z^{n}_{\tau}$ for practical purpose. The latter can be indeed readily generalized to the case of jump noise (Eqns. (99)–(100) below) without relying on stochastic calculus involving a jump measure (Section VII).
In what follows we drop the dependence of $\Phi_{n}$ on $\omega$ and sometimes on $t$, unless specified. The parameterization of the neglected scales writes then as
$\Phi_{\bm{\tau}}(X,t)=\sum_{n\geq m_{c}+1}\Phi_{n}(\tau_{n},X,t)\bm{e}_{n},\$ (47)